OpenStudy (anonymous):

Which of the following statements is false? five-ninths can be shown as equivalent to an improper fraction eight-halves can be simplified as a whole number one and one-third can be shown as equivalent to an improper fraction six−sevenths can be shown as equivalent to a proper fraction

OpenStudy (anonymous):

five-ninths can be shown as equivalent to an improper fraction becaues u cant make 5/9 improper

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Anytime and welcome to openstudy:D

OpenStudy (anonymous):

How could you eliminate the fractions to make solving the equation five-twelfths m plus two-twelfths equals eleven-twelfths easier? Divide both sides of the equation by five-twelfths. Divide both sides of the equation by twelve. Multiply both sides of the equation by twelve. Subtract two-twelfths from both sides of the equation.

OpenStudy (anonymous):

can you help me with this one?

OpenStudy (anonymous):

So 5/12m+2/12=11/12?

OpenStudy (anonymous):

Multiply both sides of the equation by twelve. because then u end up with 5m+2=11

OpenStudy (anonymous):

Which shows fifteen-tenths as a simplified improper fraction? three-halves one and one-half five-thirds one and five-tenths

OpenStudy (anonymous):

three-halves 15/10 /3 /3 3/2

OpenStudy (anonymous):

Sorry divide by 5

OpenStudy (anonymous):

thank you, one more thing

OpenStudy (anonymous):

three-fourths z minus one-sixth equals negative one-half for z. z equals negative four z equals negative eight-ninths z equals negative four-ninths z equals negative one-fourth

OpenStudy (anonymous):

z equals negative four-ninths Sorry took so long thought both sides had Z.

OpenStudy (anonymous):

It's ok

OpenStudy (anonymous):

Thought it was 3/4z -1/6--1/2z U didnt have the part "solve for Z" but that dont matter cuz we still got the answer:D

OpenStudy (anonymous):

Thank you very much. :)

OpenStudy (anonymous):

:D